Argonne National Laboratory

Advanced Photon Source
Sector 7

Argonne Home > Advanced Photon Source >

MHATT-CAT Useful Operational Data for Run 2 of FY02


Eric Dufresne, Dohn Arms, Don Walko
MHATT-CAT
Started January 21, 2002
(www.mhatt.aps.anl.gov/Sectors/Sector7/Operations/FY02_run2/)



Table of Content:

Introduction.

The work shown below is an account of useful operational data taken during
run 2 of FY02 on 7ID. This page will contain useful stability information on the
7ID High Heat Load Monochromator and also on the various repairs and problems 
identified during the run.

Summary of changes made on 7ID during the January 2002 shutdown.

During the January 2002 shutdown, the 7ID monochromator was warmed up to room
temperature in order to perform the Cryocooler yearly maintenance (bearing 
change). The bearings need to be changed every 5000 hours, but the pump can run
for 10000 hours without damage. The company says 5000 hours at 90 Hz, so we
typically run at 30 Hz so in principle the bearing life decreases as the speed
increase (linear  approximation), thus one might expect our bearings to last 
15000 Hours. I've typically changed the bearings once a year or about every 8000
to 10000 hours. 

On January 21 2002, Dohn Arms and Eric Dufresne realigned the Chi angle of 
the first Mirror Filter tank. This angle was 1.5 degree and was brought back
to 0.1 degree after removing flanges and reassembling them. This fixes an 
installation mistake uncovered after the installation in July 2000. This should
help us to align the first mirror filter more easily and allow to switch 
coating without significant realignment of MF1.

Fig. 01-21.1 shows the chi angle measured before the repair (1.5 degrees).
To realign the top of the MF1 tank, several flanges had to be redone. Fig. 01-21.2
shows one of the flanges undone. We also decided to replace the gate valve 
because we suspect it may be faulty. When the monochromator is not pumped on, we
have found that the ion pump on the MF1 trips, and during the last shutdown, 
Eric noticed that once the monochromator was vented to air, the pressure in the 
MF1 tank rose steadily to trip the ion pump. It is possible that the problem is
caused by a faulty MF1 ion pump controller too, but if we find any improvements,
then at least something will have been done to improve the uncontrolled GV3 
closing we've experienced last run during run 1 of FY02. We had thought the
problem had been fixed by removing the AGM tilt sensor but this is not the case.

Fig. 01-21.3 shows the repair nearly completed. Note that levelling Chi
also untwisted the below mounted on the monochromator and the below mounted just
before MF1. The repair thus attempted to fix several problems: chi, a periodic 
vacuum fault, and too much stress on a few bellows.  Fig. 01-21.4 shows
the final Chi angle of the first mirror tank. The angle is now 0.1 degree. It
appears that the vacuum was not affected and that no major problems developped
due to this repair. 

On January 22, Eric was able to cool the monochromator without any incidents.
The base pressure is now possibly a little higher than it was last November,
but it is still 1.7E-8 Torr with beam. Dohn Arms repaired our
7ID-B vacuum problem (see Fig. 01-22.1). The old window from 7ID-C was installed
at the end of 7ID-B and raised with an offset flange. The broken 7ID-B window
will be sent for repairs soon.  Dohn also installed an additional pumping port 
in 7ID-C to pump the P5 and the beam transport without having to reopen the P5
minihutch. Reopening always has to involve a Health Physics survey, which is a
nuisance. Fig. 01-22.2 and Fig. 01-22.3 show Dohn at work on the new port.
The P5 and Be window problem was described in pictures last year here.
MF1 chi angle before repair.

Fig. 01-21.1. Original alignment of MF1 Chi (off by 1.5 degree).

GV3 valve seal undone.

Fig. 01-21.2. Undoing flanges to align MF1 chi. We switched GV3, the gate valve shown here
with a spare.

View of nearly completed repairs.

Fig. 01-21.3. Nearly completed repairs.

Final chi angle of MF1.

Fig. 01-21.4. The final Chi angle of the top flange of MF1 (0.1 degree).

Repaired Be window in 7ID-B.

Fig. 01-22.1. The repaired Be window in 7ID-B. We used the old Be window in 7ID-C and made an offset flange to lift it up.

Dohn hard at work in 7ID-C.

Fig. 01-22.2. Dohn Arms added a new pumping port in 7ID-C so that if a
leak develops like in Nov 2001, then the P5 minihutch does not need to be reopened.

Repumping the beam transport and P5 after the Be window repair.

Fig. 01-22.3. Repumping the beam transport and P5 after the Be window repair.






Thursday, Jan. 24, 2002. Time series with a 0.5 mm (H) by 0.5 mm (V)
white beam, with the mono at 10.0 keV.


The data was started on 01/24 at 00h23 and lasts for more than 14.5 hours. Two
figures are shown below. Fig. 01-24.1 shows a time 
series of the intensity not normalized to beam current decay. the data was taken
every 3 seconds.  The diode signal tracks the ring current decay as expected.
One glitch occur in the data because of access to 7ID-B after 10 hours in the
time series. The LN2 fills occur every 2 hours or so. The 7ID-B and C ion
chamber were disabled for this tseries. 

Fig. 01-24.2 shows a time serie of the 7ID-C Beam 
Position. The beam motion peak to peak is about 110 and 50 microns in the
vertical and horizontal direction respectively. 
Intensity data started on 01/24/02.

Fig 01-24.1 . Time series of the beam intensity in 7ID-C, starting at 00h23 on 01/24
and lasting 14.5 hours. The 7ID-C diode sum, the ring current decay, and LN2 level
are the only valid data because both ion chamber in B and C were disabled.

XBPM data started on 01/24/02.

Fig 01-24.2 . The beam position, 49 m from the source, or 19 m from the High Heat Load mono
in 7ID-C during the same time series as Fig. 01-24.1. The time series is started on 01/24 at 00h23,
and lasts for 14.5 hours. The beam is stable to about 110 and 50 microns in the vertical and horizontal
directions respectively, likely due to thermal drifts.

Wednesday, Mar. 20, 2002. Beamline vacuum investigations and major EPS
water flow sensor failure.

Today, I spent the morning doing maintenance on the cryocooler vacuum jackets 
and I investigated the EPS Vacuum situation. I first outgassed all the ion 
gauges, hoping it may prevent pressure spikes, particularly on the in-vacuum
Huber chamber. 

One problem the MHATT-CAT staff should fix is to slow down the reponse of the 
goniometer ion gauge controller, the Granville Phillips 307. It can be set to 
respond in 3s rather than the default 0.5s response. It is well known that 
when the Huber circle suddenly moves, the motor shaft seal leaks and often 
will close the FEV and trip the beamline EPS. This occur when the mono has not 
been changed energy for several weeks. The beamline staff should repair this
problem before II's join in.

The other less known problem is that the L5-20 controller display may sometime 
with beam display 0 instead of the actual pressure. This is beam related, and
only occurs when X-rays are in 7ID-A. It is believed that the solution is the 
shield the outside of the ion gauge from the white beam.

At night, I investigated the X-ray BPM and now suspect a broken foil. Around 
midnight, another Proteus water flow sensor broke, but for the first time on
the L5-23 fixed mask. The beamline vacuum screen said mask flow trip. So I 
swapped the electronics module again with our spare and the unit now reads again
2.5 gpm. It took a while to reste the beamline because the APS Front End Photon
Shutter 1 (PS1) was closed and needed to be reset. Note that this is the second
such failure this month. Dohn Arms swapped the L5-20 Proteus flow sensor on
3/14/2002. This flow sensor is brand new since it replaced the broken flow
sensor we repaired in Sept. 2001. Why did it flake out so fast? Could it have
something to do with the fact that in the past two runs, we've closed the gap 
more than in the last several years, in particular pushing the mono to higher 
energies such as 20-25 keV?

Thursday, Mar. 21, 2002. Repair of the X-ray BPM and
mono energy calibration near the Zn edge.

Today, I recalibrated the mono energy with the Zn edge. The absorption 
coefficient is shown in Fig. 03-21.1. This time the scans were performed
with 0.5 eV steps, consistent with the EXAFS data taken in the elemental foil
set. Fig. 03-21.2 shows the derivative of the previous scan. The edge 
had moved by 2 eV from 9.659 to 9.661 keV.  This change may have occured
following the realignment from white beam operations to monochromatic beam
operation.

Next, the X-ray beam position monitor in 7ID-C was repaired. During the UofM 
Geology experiment, the fluorescence signal was lost. The diodes were tested 
and were still working. I also tried to read the signal from one diode with a 
SRS 570 without  success. The problem seemed to point to a broken foil. 

Don Walko and I opened the X-ray BPM vacuum T and found indeed that the foil
was shattered (see Fig. 03-21.3). We had one last spare, thus replaced it with 
a wrinkly foil (see Fig. 03-21.4). I retested it tonight and it works again!
The near edge absorption of a Zn foil.

Fig. 03-21.1. The near edge absorption coefficient of a Zn foil performed on 3/21/2002.
The edge closely resembles EXAFS data provided by the foil set vendor.

The derivative spectrum of the Zn absorption edge.

Fig. 03-21.2. The derivative spectrum of the Zn absorption edge performed on 3/21/2002.
In the foil cabinet, the edges (here 9.659 keV) is defined as the peak of the derivative spectrum.
The calibration was off by 2 eV.

Broken Cr foil in X-ray BPM in 7ID-C.

Fig. 03-21.3. The broken Cr foil inside the 7ID-C X-ray BPM was repaired on 3/21/2002
by Don Walko and Eric Dufresne. They also replaced the Be window with a thicker one (0.01").

New Cr foil inside X-ray BPM in 7ID-C.

Fig. 03-21.4. The new Cr foil inside the 7ID-C X-ray BPM. Note the wrinkles.


Back to my homepage (retour vers ma page maison).
Récemment modifié le 04/08/2002
par Eric Dufresne,
Bldg 432 Rm D007 630-252-0274
U.S. Department of Energy UChicago Argonne LLC Office of Science - Department of EnergyOffice of Basic Energy Sciences - Department of Energy
Privacy & Security Notice | Contact Us | Site Map